# Asymptotics for partitions over the Fibonacci numbers and related sequences

### v2

@@ 3. Fibonacci partitions $p_{F}(n)$ via the saddle point method
The proof of Theorem 4 was accomplished by iteratively applying an asymptotic result on exponential integrals, which can be found in the book of Olver [9, Theorem 7.1]. Note that Coons and Kirsten [3] use the notation $t_{\Lambda}^{0}(n)$ for our definition of $p_{\lambda}(n)$ above. $\Diamond$
To apply this result, we must first determine the saddle point for large values of $n$. From (13), we have that
@@
For sufficiently large $n$, or, equivalently, for sufficiently small $\alpha>0$,
$n=\frac{-\log(\alpha)}{\alpha\log\varphi}\left(1+O\left(\frac{1}{\log\alpha}\right)\right).$
@@
We calculate
@@
$=-\left.\frac{d}{ds}\left(\frac{(\log s)^{2}}{2\log\varphi}-(\log s)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(s)+O(s)\right)\right|_{s=\alpha}$
(15) $=-\frac{\log\alpha}{\alpha\log\varphi}+O\left(\frac{1}{\alpha}\right),$
which, after factoring out the first two terms gives the result. ∎
@@
$\alpha=\frac{1}{n\log\varphi}\left(1+O\left(\frac{1}{\log n}\right)\right),$
and
$\left.\frac{dn}{ds}\right|_{s=\alpha}=-n^{2}(\log n)\log\varphi\left(1+O\left(\frac{1}{\log n}\right)\right).$
@@
It is quite clear that $O(\alpha/\log\alpha)=O(1/n).$ For the rest, we note that if $\log x=y\log y$, then for large $x$, $y=W(x)$, where $W$ is Lambert’s $W$-function, which satisfies
$W(x)=\log x(1+O(\log\log x/\log x))$
for all $x$ sufficiently large. Here, we have $x=\exp(n\log\varphi)$ and $y=1/\alpha$, so that
$\alpha=\frac{1}{W(\exp(n\log\varphi))}\left(1+O\left(\frac{1}{\log n}\right)\right)=\frac{1}{n\log\varphi}\left(1+O\left(\frac{1}{\log n}\right)\right),$
which proves the first assertion.
For the second assertion, we use (15) together with the first assertion to give
$\left.\frac{dn}{ds}\right|_{s=\alpha}=\frac{\log\alpha}{\alpha^{2}\log\varphi}\left(1+O\left(\frac{1}{\log\alpha}\right)\right)=-n^{2}(\log n)\log\varphi\left(1+O\left(\frac{1}{\log n}\right)\right).\qed$
We have the following immediate corollary of the first assertion of Proposition 6.
@@
$\log\alpha=-\log n-\log\log\varphi+O\left(\frac{1}{\log n}\right).$
@@
$f(\alpha)=\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n^{2}}\right),$
for some $1$-periodic function $\psi(x)$, and any $\varepsilon>0$.
@@
As $s\to 0^{+}$, using Proposition 4 and one application of L’Hôpital’s rule gives $g(s)=O(s^{2})$ (in fact, one gets that $g(s)=O(s^{4-\varepsilon})$ for any fixed small positive $\varepsilon$, but only $O(s^{2})$ is necessary after later comparison with the asymptotics for $h(s)$). For the result on $f(s)$, note that collectively, the sum of all of the terms with $k>0$ in the formula for $h(s)$ in Proposition 5 go to zero as $s\to 0^{+}$, since $\Re(-s(n,k))=2k>0$. Thus, as $s=\alpha\to 0^{+}$, separating out the $k=0$ term of the sum, and noting that $s(n,0)=\frac{2\pi in}{\log\varphi}$, we have,
(18) $f(\alpha)$ $=\frac{1}{\log\varphi}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}\left(\frac{\alpha}{\sqrt{5}}\right)^{-\frac{2\pi in}{\log\varphi}}\Gamma(\tfrac{2\pi in}{\log\varphi})\,\zeta(1+\tfrac{2\pi in}{\log\varphi})+O(\alpha^{2})$
@@
Since both the sine and cosine functions are $2\pi$-periodic, (19) gives that $f(s)=f(\varphi s)+O(s^{2})$ as $s\to 0^{+}$. Applying Proposition 6, we obtain
$\frac{2\pi j}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)=-2\pi j\left(\frac{\log n}{{\log\varphi}}+\frac{\log\log\varphi+\log\sqrt{5}}{\log\varphi}\right)+O\left(\frac{1}{\log n}\right).$
Finally, we have $O(\alpha^{2})=O(n^{-2})$, which finishes the proof. ∎
@@
Proof of Theorem 1 for $p_{F}(n)$.
We evaluate the pieces of the asymptotic in Theorem 4. Towards this end, the two parts of Proposition 6 give, respectively,
(20) $e^{n\alpha}=e^{n\cdot\frac{1}{n\log\varphi}\left(1+O\left(\frac{1}{\log n}\right)\right)}=e^{\frac{1}{\log\varphi}}\left(1+O\left(\frac{1}{\log n}\right)\right)$
and
(21) $\left(\sqrt{\frac{1}{-\left.\frac{dn}{ds}\right|_{s=\alpha}}}+O\left(\frac{1}{n^{3/2}}\right)\right)=\frac{1}{n}\left(\log n\log\varphi\right)^{-1/2}\left(1+O\left(\frac{1}{\log n}\right)\right).$
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $F_{2}(e^{-\alpha})$. To this end, we combine Theorem 3 with Corollary 1 and Proposition 7 to obtain, as $\alpha\to 0^{+}$,
$\log F_{2}(e^{-\alpha})$ $=\frac{(\log\alpha)^{2}}{2\log\varphi}-(\log\alpha)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(\alpha)+O(\alpha)$
$=\frac{1}{2\log\varphi}(\log n)^{2}\left(1+O\left(\frac{1}{(\log n)^{2}}\right)\right)$
$\quad+\left(\log n+\log\log\varphi+O\left(\frac{1}{\log n}\right)\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n}\right)$
$=\frac{(\log n)^{2}}{2\log\varphi}\left(1+O\left(\frac{1}{(\log n)^{2}}\right)\right)+(\log n)\left(\frac{c_{3}}{\log\varphi}-1\right)$
$\quad+\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{(\log n)^{2}}\right),$
where $\psi(x)$ is the $1$-periodic function described in Proposition 7. Thus, we have
$F_{2}(e^{-\alpha})=\psi_{1}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log n}{2\log\varphi}\left(1+O\left(\frac{1}{(\log n)^{2}}\right)\right)+\frac{c_{3}}{\log\varphi}-1\right)}\left(1+O\left(\frac{(\log n)^{2}}{n}\right)\right),$
where $\psi_{1}(x)$ is the $1$-periodic function satisfying
$\psi_{1}\left(\frac{\log n}{\log\varphi}\right):=\exp\left(\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)\right).$
Combing the asymptotic for $F_{2}(e^{-\alpha})$ with equations (20) and (21) gives the desired result, with
$\psi_{p_{F}}(x):=\frac{e^{\frac{1}{\log\varphi}}\psi_{1}(x)}{\sqrt{2\pi\log\varphi}}.\qed$

### v3

@@ 3. Fibonacci partitions $p_{F}(n)$ via the saddle point method
The proof of Theorem 4 was accomplished by iteratively applying an asymptotic result on exponential integrals, which can be found in the book of Olver [11, p. 127, Theorem 7.1]. Note that Coons and Kirsten [3] use the notation $t_{\Lambda}^{0}(n)$ for our definition of $p_{\lambda}(n)$ above. $\Diamond$
Remark 2.
The statement “$\alpha$ must be thought of as being replaced by its large-$n$ asymptotic expansion” may seem a bit cumbersome, but, here, the point is that the solution of (12) gives $n$ as a function of the saddle point $\alpha$ as an (asymptotically) monotonic function, so it is invertible; that is, there is a well-defined asymptotic for the saddle point $\alpha$ in terms $n$—this is precisely the method we employ.
To apply this Theorem 4, we must first determine the saddle point for large values of $n$. From (12), we have that
@@
There exists a function $h_{0}(s)$ satisfying $h_{0}(s)=h_{0}(\varphi s)$ such that for sufficiently large $n$, or, equivalently, for sufficiently small $\alpha>0$,
$n=\frac{-\log(\alpha)}{\alpha\log\varphi}+\frac{h_{0}(\alpha)}{\alpha}+O\left(\alpha\right).$
@@
Let $f(s):=g(s)+h(s)$, where $g(s)$ and $h(s)$ are as in Propositions 4 and 5, respectively. As $s\to 0^{+}$, using Proposition 4 and one application of L’Hôpital’s rule gives $g(s)=O(s^{2})$ (in fact, one gets that $g(s)=O(s^{4-\varepsilon})$ for any fixed small positive $\varepsilon$, but only $O(s^{2})$ is necessary after later comparison with the asymptotics for $h(s)$). Now, collectively, the sum of all of the terms with $k>0$ in the formula for $h(s)$ in Proposition 5 go to zero as $s\to 0^{+}$, since $\Re(-s(n,k))=2k>0$. Thus, as $s=\alpha\to 0^{+}$, separating out the $k=0$ term of the sum, denoting it $[k=0]h(s)$, and noting that $s(n,0)=\frac{2\pi in}{\log\varphi}$, we have,
$f(\alpha)=[k=0]h(\alpha)+O(\alpha^{2})=\frac{1}{\log\varphi}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}\left(\frac{\alpha}{\sqrt{5}}\right)^{-\frac{2\pi in}{\log\varphi}}\Gamma(\tfrac{2\pi in}{\log\varphi})\,\zeta(1+\tfrac{2\pi in}{\log\varphi})+O(\alpha^{2}),$
so that
$f^{\prime}(\alpha)=\frac{\sqrt{5}}{\alpha}\cdot[k=0]h(\alpha)+O(\alpha).$
The form of $[k=0]h(s)$ immediately implies that $[k=0]h(s)=[k=0]h(\varphi s)$.
Now, we calculate
@@
$=-\left.\frac{d}{ds}\left(\frac{(\log s)^{2}}{2\log\varphi}-(\log s)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(s)+O(s^{2})\right)\right|_{s=\alpha}$
(14) $=-\frac{\log\alpha}{\alpha\log\varphi}-\frac{1}{\alpha}\left(\frac{c_{3}}{\log\varphi}-1-\sqrt{5}\cdot[k=0]h(\alpha)\right)+O\left(\alpha\right).$
setting $h_{0}(s):=\sqrt{5}\cdot[k=0]h(s)+1-c_{3}/\log\varphi$ gives the result. ∎
Remark 3.
Lemma 3 provides the leading two terms for $n$ in terms of the saddle point $\alpha$. Here, the first term shows that the relationship is asymptotically monotonic, so that it can be inverted. The second term is oscillatory. If one follows this method and tries to apply it to the distinct Fibonacci partitions function $q_{F}(n)$ the resulting asymptotic is not monotonic—the leading term is oscillatory. This is precisely why this method doesn’t immediately generalise to distinct Fibonacci partitions.
While the above lemma gives $n$ as a function of $\alpha$, we necessarily need $\alpha$ as a function of $n$ to apply the saddle point method. We achieve this via the Lambert $W$-function.
@@
There is a continuous $1$-periodic function $\psi_{0}(x)$, such that for sufficiently small $\alpha>0$, or, equivalently, for sufficiently large $n$,
$\alpha=\frac{W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)}{n\log\varphi}\left(1+O\left(\frac{\log n}{n^{2}}\right)\right),$
where $W(x)$ denotes Lambert’s $W$-function.
@@
Note that the previous lemma gives that
$n=\left(\frac{-\log(\alpha)}{\alpha\log\varphi}+\frac{h_{0}(\alpha)}{\alpha}\right)\left(1+O\left(\frac{\alpha^{2}}{\log{\alpha}}\right)\right),$
where $h_{0}(s)$ is fixed along any sequence $\{x\varphi^{m}\}_{m\geqslant 0}$. We use this property to invert the above relationship between $\alpha$ and $n$. Note that the relationship is invertible because the lead asymptotics are strictly monotonic. Now, when one inverts $n\sim A\frac{1}{\alpha}\log\frac{1}{\alpha}+B\frac{1}{\alpha}$, one gets $\alpha\sim A\cdot W\left(e^{B/A}\frac{n}{A}\right)/n\sim A\log n/n$, where $W(x)$ is Lambert’s $W$-function. Doing this along the sequences $\{x\varphi^{m}\}_{m\geqslant 0}$ to ensure a constant $B=B(x)$, we then reconstruct, using a fundamental interval, say $x\in[\varphi,\varphi^{2}]$, to get a continuous $1$-periodic function $\psi_{0}(x)$ such that for large $n$,
$\alpha=\frac{W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)}{n\log\varphi}\left(1+O\left(\frac{\log n}{n^{2}}\right)\right).$
Here, we have used the original relationship to find that $O(\alpha/\log\alpha)=O(1/n),$ and then the inverse, noting that $W(n)\sim\log n$. ∎
In what follows, we will use Proposition 6 to give asymptotics for several functions of $\alpha$, including $n\alpha$, $\left.\frac{dn}{ds}\right|_{s=\alpha}$, $\log\alpha$ and $(\log\alpha)^{2}$. To give our end result, we will need varying orders of precision for the asymptotics for each of these terms. In particular, first order asymptotics of the Lambert $W$-function will not be enough to deal with $(\log\alpha)^{2}$, though they will be enough for some terms so we record them below. For convenience, we note here that the Lambert $W$-function satisfies,
(15) $W(x)=\log x-\log\log x+\frac{\log\log x}{\log x}+O\left(\frac{\log\log x}{(\log x)^{2}}\right),$
as $x\to\infty$; see, e.g., Corless et al. [5].
@@
$\alpha=\frac{\log n}{n\log\varphi}\left(1+O\left(\frac{\log\log n}{\log n}\right)\right),$
and
$\log\alpha=-\log n+\log\log n-\log\log\varphi+O\left(\frac{\log\log n}{\log n}\right).$
Proof.
The first result follows directly from the fact that
$W(x)=\log x(1+O(\log\log x/\log x))$
for all $x$ sufficiently large. The second follows immediately from the first. ∎
Corollary 2.
$\left.\frac{dn}{ds}\right|_{s=\alpha}=\frac{-n^{2}\log\varphi}{\log n}\left(1+O\left(\frac{\log\log n}{\log n}\right)\right).$
Proof.
We start with Lemma 3 in the form $n=-\log(\alpha)/(\alpha\log\varphi)+O\left({1}/{\alpha}\right),$ and take a derivative, then apply both parts of Corollary 1 to get
$\left.\frac{dn}{ds}\right|_{s=\alpha}=\frac{\log\alpha}{\alpha^{2}\log\varphi}+O\left(\frac{1}{\alpha^{2}}\right)=\frac{\log\alpha}{\alpha^{2}\log\varphi}\left(1+O\left(\frac{1}{\log\alpha}\right)\right)=\frac{-n^{2}\log\varphi}{\log n}\left(1+O\left(\frac{\log\log n}{\log n}\right)\right),$
which finishes the proof. ∎
The final term necessary is $(\log\alpha)^{2}$. Here, in addition to using the full asymptotic in (15), we will use the fact that for any $y$, $\log W(y)=\log y-W(y)$.
Corollary 3.
$(\log\alpha)^{2}=\log n\left(\log n-2\log\log n-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi\right)\\ +\log\log n\left(\log\log n+2-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi\right)\\ +(\log\log\varphi)^{2}-\frac{2\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi+O\left(\frac{(\log\log n)^{2}}{\log n}\right).$
Proof.
We start with Proposition 6 and take the natural logarithm of both sides to obtain
(16) $(\log\alpha)^{2}=\left(\log\left(\frac{W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)}{n\log\varphi}\right)\right)^{2}+O\left(\frac{(\log n)^{2}}{n^{2}}\right),$
since
$\log\left(\frac{W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)}{n\log\varphi}\right)=O(\log n).$
We now use the full force of (15) to give an asymptotic for the first term in (16), first noting that
$\log\left(\frac{W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)}{n\log\varphi}\right)$ $=\log\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)$
$\qquad\qquad-\log(n\log\varphi)-W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)$
$=\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}-W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right).$
Thus,
(17) $\left(\log\left(\frac{W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)}{n\log\varphi}\right)\right)^{2}=\left(\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}\right)^{2}\\ -\frac{2\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)+W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)^{2}.$
For the middle term we use the first three terms of the asymptotic of $W$ in (15) and, for the square, we will use the square of all of (15), which is, as $x\to\infty$,
$W(x)^{2}=(\log x)^{2}+(\log\log x)^{2}-2\log x\log\log x+2\log\log x+O\left(\frac{(\log\log x)^{2}}{\log x}\right).$
To this end, we to determine strong estimates for the asymptotics of $\log x$, $\log\log x$ and $\log x\log\log x$ with $x=e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi.$ Here, we have
(18) $\log\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)$ $=\log n+\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}-\log\log\varphi$
$=\log n\left(1+\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi-\log\log\varphi}{\log n}\right),$
and so, using that as $y\to 0$, $\log(1+y)=y-y^{2}/2+O(y^{3})$, we have
(19) $\log\log\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)=\log\log n+\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi-\log\log\varphi}{\log n}+O\left(\frac{1}{(\log n)^{2}}\right).$
So, using the above asymptotics,
(20) $W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)=\log n+\log\log n+\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}-\log\log\varphi+O\left(\frac{\log\log n}{\log n}\right),$
and
(21) $W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)^{2}=(\log n)^{2}-2\log n\log\log n\\ -2\log n\left(\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}-\log\log\varphi\right)+(\log\log n)^{2}+2\log\log n\left(1+\log\log\varphi-\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}\right)\\ +\left(\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}-\log\log\varphi\right)^{2}-2\left(\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}-\log\log\varphi\right)+O\left(\frac{(\log\log n)^{2}}{\log n}\right).$
Combining (16), (17), (20) and (21) gives the result. ∎
@@
$f(\alpha)=\psi_{1}\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{\log\log n}{\log n}\right),$
for some $1$-periodic function $\psi_{1}(x)$, and any $\varepsilon>0$.
@@
We proceed as in the proof of Lemma 3, and, as in that proof, separating out the $k=0$ term of the sum for $f(\alpha)$ and denoting it $[k=0]h(s)$, as $s=\alpha\to 0^{+}$,
$f(\alpha)=[k=0]h(\alpha)+O(\alpha^{2})$ $=\frac{1}{\log\varphi}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}\left(\frac{\alpha}{\sqrt{5}}\right)^{-\frac{2\pi in}{\log\varphi}}\Gamma(\tfrac{2\pi in}{\log\varphi})\,\zeta(1+\tfrac{2\pi in}{\log\varphi})+O(\alpha^{2})$
@@
Since both the sine and cosine functions are $2\pi$-periodic, (22) gives that $f(s)=f(\varphi s)+O(s^{2})$ as $s\to 0^{+}$. Applying Corollary 1, we obtain
$\frac{2\pi j}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)=-2\pi j\left(\frac{\log n}{{\log\varphi}}\right)+O\left(\frac{\log\log n}{\log n}\right).$
Finally, we note that $O(\alpha^{2})=O((\log n)^{2}/n^{2})=O\left(\log\log n/\log n\right)$ and set $\psi_{1}(s)=[k=0]h(s)$ to finish the proof. ∎
@@
Proof of Theorem 1.
We evaluate the pieces of the asymptotic in Theorem 4. Towards this end, Proposition 6 combined with (20) gives,
${n\alpha}$ $={\frac{1}{\log\varphi}W\left(e^{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)/\log\varphi}n/\log\varphi\right)\left(1+O\left(\frac{\log n}{n^{2}}\right)\right)}$
$=\frac{\log n}{\log\varphi}+\frac{\log\log n}{\log\varphi}+\frac{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{(\log\varphi)^{2}}-\frac{\log\log\varphi}{\log\varphi}+O\left(\frac{\log\log n}{\log n}\right),$
so that
(23) $e^{n\alpha}=e^{\frac{1}{(\log\varphi)^{2}}{\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}-\frac{\log\log\varphi}{\log\varphi}}n^{1/\log\varphi}(\log n)^{1/\log\varphi}\left(1+O\left(\frac{\log\log n}{\log n}\right)\right)$
Corollary 2 gives
(24) $\left(\sqrt{\frac{1}{-\left.\frac{dn}{ds}\right|_{s=\alpha}}}+O\left(\frac{1}{n^{3/2}}\right)\right)=\frac{1}{n}\left(\frac{\log n}{\log\varphi}\right)^{1/2}\left(1+O\left(\frac{\log\log n}{\log n}\right)\right).$
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $F_{2}(e^{-\alpha})$. To this end, we combine Theorem 3 with Corollaries 1 and 3, and Proposition 7 to obtain, as $n\to\infty$,
$\log F_{2}(e^{-\alpha})$ $=\frac{(\log\alpha)^{2}}{2\log\varphi}-(\log\alpha)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(\alpha)+O(\alpha^{2})$
$=\frac{\log n}{2\log\varphi}\left(\log n-2\log\log n-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi+2c_{3}-2\log\varphi\right)$
$\qquad+\frac{\log\log n}{2\log\varphi}\left(\log\log n+2-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi-2c_{3}+2\log\varphi\right)$
$\qquad\qquad+\frac{(\log\log\varphi)^{2}}{2\log\varphi}-\frac{2\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{2(\log\varphi)^{2}}+\frac{\log\log\varphi}{\log\varphi}+\log\log\varphi\left(\frac{c_{3}}{\log\varphi}-1\right)$
$\qquad\qquad\qquad+c_{2}+2\gamma+\psi_{1}\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{(\log\log n)^{2}}{\log n}\right).$
Thus, we have
$F_{2}(e^{-\alpha})=\psi_{2}(n)n^{a(n)}(\log n)^{b(n)}\left(1+O\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
where $\psi_{2}(n)$ is the strictly positive bounded (above and below) function
$\psi_{2}\left(n\right)$ $:=\exp\Bigg{(}\frac{(\log\log\varphi)^{2}}{2\log\varphi}-\frac{2\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{2(\log\varphi)^{2}}+\frac{\log\log\varphi}{\log\varphi}$
$\qquad\qquad\qquad+\log\log\varphi\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi_{1}\left(\frac{\log n}{\log\varphi}\right)\Bigg{)},$
$a(n)$ $:=\frac{1}{2\log\varphi}\left(\log n-2\log\log n-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi+2c_{3}-2\log\varphi\right)$
$b(n)$ $:=\frac{1}{2\log\varphi}\left(\log\log n+2-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi-2c_{3}+2\log\varphi\right).$
Combing the asymptotic for $F_{2}(e^{-\alpha})$ with equations (23) and (24) gives the desired result. ∎